OpenStudy (anonymous):

Simplify. –50 + (–10).

OpenStudy (anonymous):

-60

OpenStudy (anonymous):

these are really asy, try to solve them yourself?

OpenStudy (anonymous):

lol i know but tht wouldnt be a negative because they both are negative

OpenStudy (anonymous):

Its the same as adding 50 and 10. You get -60

OpenStudy (anonymous):

-60, try to do them yourself

OpenStudy (anonymous):

it would be a negative starr, you're adding 2 negatives together. You're thinking of multiplication!

OpenStudy (anonymous):

why is that because when u have one negative how could it still be a negative with 2

OpenStudy (anonymous):

No, they only become positive's if you multiply 2 negatives.

OpenStudy (anonymous):

ya but before it was a problem like -6 +1=-7 how is it still a negative

OpenStudy (anonymous):

-6+1 is 5, because your adding a positive to neagtive --- + --- --- --

OpenStudy (anonymous):

When adding 2 negatives, or a negative and a positive, or multiplying a negative and a positive, the answer will always be a negative. :)

OpenStudy (anonymous):

It will only become positive by multiplying 2 negatives. Here are some examples: -5 + 3 = -2 -6 + -3 = -9 -4 * 2 = -8 -4 * -4 = 16